A cochlea implant includes a 10 cm2 conductor loop that is subject to time-varying magnetic field of 1 T/s, as the patient moves into the magnetic field, and a maximum magnetic field of 1.5 T in an MRI machine. The resistance of this conductor loop is 1 X 10-7 Ω. a) What is the maximum current in the conducting loop? b) What is the maximum torque exerted on the conducting loop?
Answer #1

A] Current induced = emf/Resistance = rate of change of flux/resistance = [Area*dB/dt]/R = 10e-4*1/1e-7 = 10000 A answer

B] Maximum torque = I*Area*B = 10000*10e-4*1.5 = 15 Nm